[Brahms-l] Shift report 20040119 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Mon Jan 19 2004 - 16:15:17 EST
------------------------------------------------------------
Date: Mon Jan 19 16:15:16 2004, Shift: 08:00-16:00
Supervisor: HI, with: 
Log book#15 pages 37 - 37

Shift Summary:
------------------------------------------------------------
1.  beam for 1st 2hr.  Dana and I used to debug trigger and detectors.  
2.  No data taking.
3.  T3 HV is off until alcohole bubler is fixed.  It does not hold voltage or it has high current. (See E-log).


------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jan 19 16:15:36 2004

This archive was generated by hypermail 2.1.8 : Mon Jan 19 2004 - 16:15:54 EST